# PatchTrack: A Comprehensive Analysis of ChatGPT’s Influence on Pull Request Outcomes

### v1

@@ 4.3 Discussion and Implications - RQ1
The predominance of low- and moderate-integration cases highlights key challenges in AI-assisted development, such as misalignment with project coding styles, maintainability concerns, and trust issues. This underscores the need for AI tools to incorporate better contextual awareness, generating suggestions that require minimal refinement. Although ChatGPT reduces the initial effort to write code, its effectiveness depends on the developer’s ability to evaluate, adapt, and refine AI-generated patches. At the same time, high-integration cases demonstrate that well-aligned AI-generated patches can streamline the pull request process, reinforcing AI’s potential in accelerating development workflows.
@@
In particular, our study focuses on self-admitted ChatGPT usage (SACU), which means that all the cases analyzed involve developers explicitly stating their use of ChatGPT within pull request discussions. This ensures transparency in identifying AI-assisted contributions, but also introduces a limitation: we do not capture instances where ChatGPT was used privately or without disclosure. The question of whether undisclosed ChatGPT usage follows a different pattern of patch integration, which could influence the observed distribution of patch applied (PA), patch not applied (PN), and no existing patch (NE) cases remains open.
These observations align with the longstanding vision of the “Programmer’s Apprentice” [15], which imagined AI as a reasoning collaborator in software development workflows. In this vision, AI acts not as an autonomous agent, but as a junior partner that supports design exploration, debugging, and solution refinement while leaving final decisions to the developer. Our findings reflect this paradigm: developers often treat the ChatGPT output as a starting point, adapt it based on the project-specific context, and integrate it selectively. This underscores the role of ChatGPT as a decision-support tool rather than a turn-key solution in pull request workflows.
Although our analysis quantified the extent to which ChatGPT-generated patches were integrated, it did not perform a qualitative examination of why some patches were applied more successfully than others. Future work could investigate the characteristics of applied patches, such as query formulation, patch complexity, and developer perception of AI-generated solutions. Understanding these aspects could further refine AI-assisted coding and improve AI effectiveness in collaborative development. Furthermore, this discussion sets the groundwork for RQ2 (why patches are not applied), RQ3 (why no patches are suggested), and RQ4 (why pull requests are closed despite the participation of ChatGPT), facilitating a deeper examination of the role of AI in pull request decision-making.
@@ 5 Results and Discussion – RQ2, RQ3 & RQ4
Recall that to address RQ2, RQ3, and RQ4, in Section 3.2–Step 6, we employed a rigorous qualitative analysis methodology involving summarizing ChatGPT-developer conversations and GitHub pull requests, followed by thematic extraction using a card sorting technique.
For brevity, we present one representative instance per theme in the results sections. Additional examples and detailed summaries are available in the replication package [39]. The full set of data, including all summaries, identified themes, pull request links, and corresponding ChatGPT-developer conversation links, is provided in the replication package to support reproducibility and further analysis. Each instance is labeled with a unique identifier: PN-X (X being the instance number) for patches not applied, NE-X for cases where no patches were suggested and CL-X for closed pull requests, to facilitate cross-referencing within the replication package [39].
@@
5.1 Results and Discussion - RQ2
Why are ChatGPT-suggested patches in conversations not directly integrated (PN) into the pull request, and how do developers use ChatGPT’s suggestions in their workflow under the self-admitted ChatGPT usage?
In our analysis, the tool initially classified 63 patch not applied (PN) instances. However, as detailed in Section 4.1, a manual review was performed on a sample of 55 PN cases to validate the precision of the classification. This review identified 3 false positives that were incorrectly classified as PN and 4 false negatives from patch applied (PA) that should have been included in PN. After these corrections, we obtained 56 valid PN instances for thematic analysis. The analysis of these 56 PN instances revealed six distinct themes, as illustrated in Figure 7.
@@ 5.1 Results and Discussion - RQ2
Technical limitation and preferences: This was the second most prominent theme, with 9 of 56 instances demonstrating how ChatGPT’s patches were not applied due to technical constraints or developer preferences for alternative solutions. In these cases, developers faced trade-offs in implementation strategies, either due to limitations in the existing system or subjective choices regarding the best approach to a problem. A representative example is pull request #171 [62] within the ory/elements project (PN-11), where developers attempted to disable a button when submitting the form. A reviewer sought ChatGPT’s input on a possible fix, asking: “Wouldn’t the fix just be to change click to submit? […]” However, another reviewer pointed out a critical issue with this approach: “This doesn’t actually work. The reason is that the click event fires before the submit event […] but IMO we should just get this right […]” The pull request author ultimately acknowledged that resolving the issue properly would require additional time and effort, stating: “[…] just don’t want to spend time on getting this right, right now […]” Upon further examination, we found that the proposed solution was reverted because disabling the button caused unintended side effects, such as stopping event propagation. Given the minor impact of the issue, the team decided to focus on more pressing updates rather than implementing a workaround.
@@ 5.1.1 Implications for RQ2
Our findings indicate that developers rarely apply ChatGPT-suggested patches verbatim. Instead, they critically assess AI-generated suggestions, modifying them to align with project requirements, maintainability constraints, and workflow considerations. Even when patches are not directly integrated, ChatGPT contributes to software development by shaping decision-making, guiding optimizations, and supporting documentation improvements. These insights have direct implications for both practitioners and researchers in understanding how AI-assisted workflows function when patches are not applied. Recent empirical evaluations have highlighted fundamental limitations in LLMs’ ability to understand, reason about, and modify source code [67, 68, 69]. These limitations, including difficulties in semantic comprehension and contextual adaptation, may partly explain why ChatGPT-suggested patches often required developer refinement or were ultimately not applied in their original form. Our findings reinforce the need for critical evaluation of AI-generated suggestions within collaborative development workflows.
@@ 5.2 Results & Discussion - RQ3
From the generated summaries of the 84 NE instances, the employed card sorting analysis yielded four distinct themes, as illustrated in Figure 8. The identified themes are: Conceptual Guidance & Theoretical Advice, Documentation, Communication & Translation, Education and Knowledge Sharing, and Debugging & Optimization Strategies.
@@
[…]
@@ 5.2.1 Implications for RQ3
5.3 Results & Discussion - RQ4
How does ChatGPT influence closed pull requests, and what factors contribute to these outcomes under the self-admitted ChatGPT usage? To address this research question, we conducted an in-depth analysis of 47 closed pull requests that involved interactions between developers and ChatGPT. The qualitative analysis identified six key themes, as illustrated in Figure 9. The most prevalent reason for pull request closure was attributed to Scope and Quality Issues, which accounted for 31.9% of the cases. Other reasons included procedural and technical adjustments (19.1%), experimental solutions (17.0%), administrative or policy issues (10.6%), duplicate solutions (10.6%) and cases where no specific reason (10.6%) was documented.
@@ 5.3 Results & Discussion - RQ4
Experimental Solutions: 8 of the 47 closed pull requests were later replaced by more refined implementations. These pull requests were often experimental, serving as proof-of-concept solutions that were ultimately superseded by optimized alternatives. In some cases, developers identified existing solutions that made ChatGPT-assisted contributions unnecessary for direct integration. For example, pull requests #2214 [76] and #2212 [77] in the open-learning-exchange/myplanet repository (CL-29 and CL-36) introduced ChatGPT-generated French and Arabic translations for UI strings, addressing issues #2216555https://github.com/open-learning-exchange/myplanet/issues/2216 and #2219666https://github.com/open-learning-exchange/myplanet/issues/2219. However, both pull requests were later closed, with their contributions consolidated into pull request #2229777https://github.com/open-learning-exchange/myplanet/pull/2229, which was merged. This closure pattern aligns with the common practice of refining and merging multiple contributions into a single comprehensive pull request. Although ChatGPT facilitated the initial translation drafts, final refinements were likely developer-led to ensure linguistic accuracy and adherence to project standards. This underscores how ChatGPT-assisted contributions often act as stepping stones toward more polished, project-aligned solutions rather than directly leading to merged changes.
@@
[…]
@@ 6 Related Work
Pull request decision-making: In collaborative software engineering and open-source projects, the decision-making process for pull requests is crucial to determine both the quality and the efficiency of the development. Gousios et al. [6] and Yu et al. [81] have examined how contributions are evaluated, highlighting the importance of peer review and social interactions. Various factors influence pull request acceptance, including technical merit, code quality, and alignment with project goals. For example, Tsay et al. [82] found that reviewers prioritize maintainability, readability, and adherence to coding guidelines when evaluating contributions. Similarly, Zhao et al. [20] showed that pull requests that lack comprehensive tests are more likely to be rejected, reinforcing the importance of thorough validation.
Beyond technical factors, social and ecosystem-wide dynamics also influence pull request decision-making. Dey and Mockus [22] examined how technical attributes and the reputation of contributors affect the acceptance of pull requests in the NPM ecosystem, highlighting the importance of project constraints and community participation. Storey et al. [14] further emphasized that communication dynamics across different channels profoundly shape participatory cultures in software development, underscoring the collaborative challenges that arise in distributed workflows such as pull requests. In addition to social factors, prior research has highlighted the value of mining developer disclosures to better understand development practices. Yu et al. [83] introduced Jitterbug, a machine learning-based approach for detecting self-admitted technical debt (SATD) in code comments. Inspired by this idea, our study extends the concept of self-admitted artifacts to AI-assisted development by analyzing self-admitted Chat
@@
Beyond studies on AI-generated code, prior work has investigated how automation influences developer workflows and decision-making. Cassee et al. [90, 91] showed that automated systems such as Continuous Integration (CI) and GitHub bots alter code review behaviors and are critically evaluated by developers based on trust and perceived utility. Extending this line of inquiry, Shihab et al. [92, 93] analyzed the adoption of LLM-based chatbots and automation bots in pull request workflows, emphasizing challenges related to trust, accuracy, and context sensitivity. These findings underscore the need to examine how developers interact with suggested ChatGPT patches, critically assess them, and integrate or modify AI-assisted contributions during collaborative development.
In addition to studies on AI-generated code retention and quality, researchers have broadly examined AI-assisted software development, including the roles of ChatGPT and GitHub Copilot in code reviews, debugging, and software documentation [27, 28, 29, 94, 95, 96, 9]. Although these studies demonstrate improvements in productivity and code quality, they often focus on controlled experiments rather than real-world open-source development [30, 31, 32]. Previous work has explored the role of AI in pull request workflow automation [18, 35], but little research has examined how developers interact with ChatGPT in pull request discussions and adapt AI-generated patches to fit project-specific constraints. Our work addresses this gap by empirically analyzing ChatGPT’s impact in pull request workflows, particularly the decision-making processes that influence patch acceptance, modification, or rejection.

### v2

@@ 4.2.3 Discussion and Implications – RQ1
The predominance of low- and moderate-integration cases underscores ongoing challenges in AI-assisted development, including misalignment with project coding styles, maintainability concerns, and trust calibration. This suggests a need for AI tools that incorporate greater contextual awareness, generating code that requires less reshaping. At the same time, high-integration cases demonstrate AI’s potential to streamline the pull request process when outputs align with project expectations, reducing the effort of implementation.
@@
Importantly, our study focuses on self-admitted ChatGPT usage (SACU), which means that all the cases analyzed involve developers explicitly stating their use of ChatGPT within pull request discussions. This ensures transparency in identifying AI-assisted contributions, but also introduces a limitation: we do not capture instances where ChatGPT was used privately or without disclosure. The question of whether undisclosed ChatGPT usage follows a different pattern of patch integration, which could influence the observed distribution of PA, PN, and NE cases remains open.
These findings resonate with the longstanding vision of the “Programmer’s Apprentice” [21], positioning ChatGPT not as an autonomous coder but as a junior collaborator that supports debugging, scaffolding, and refinement while leaving final decisions to developers. By classifying adoption behaviors across granular integration levels, we show that ChatGPT functions primarily as a decision-support tool, shaping but not replacing human judgment in pull request workflows.
Finally, this analysis establishes a foundation for the subsequent research questions. While RQ1 quantifies how ChatGPT-generated patches are integrated and characterizes adoption patterns, RQ2–RQ4 explore the complementary cases: why patches are not applied (PN), why no patches are suggested (NE), and why some pull requests close despite ChatGPT’s involvement (CL).
Connection to Prior Work: Our findings complement and extend prior studies on AI-assisted code generation [3, 1, 6], code review [36, 20], and pull request workflows [15, 16, 17]. While prior work has quantified productivity benefits or patch acceptance [29, 30, 31], few studies have examined how AI-generated patches are actually evaluated, reshaped, or rejected within the collaborative dynamics of pull requests. Our contribution builds on recent efforts to capture patch-level decisions [10, 11], but goes further by stratifying cases across integration quartiles and classifying behavioral themes. In doing so, we show that generative AI functions less as a deterministic patch source and more as a catalyst for developer reasoning and selective adoption.
@@ 5 Results and Discussion – RQ2, RQ3 & RQ4
This section builds on the validated dataset from RQ1, where 230 pull requests were manually reviewed to assess patch integration outcomes: 90 PA, 56 PN, and 84 NE. We now examine the PN and NE subsets to answer RQ2 and RQ3. For RQ4, we independently sampled 47 closed pull requests from a total of 53 using a 95% confidence level and a 5% margin of error.
Thematic coding was conducted using a structured framework based on our research questions. While we did not compute Cohen’s Kappa due to the interpretive nature of the task, agreement was reached through coder discussion and convergence on final themes. Details of the coding procedure and reviewer roles are described in Section 3.2, Step 5.
For brevity, we present one representative instance per theme in the results sections. Additional examples and detailed summaries are available in the replication package [42], which includes all summaries, identified themes, pull request links, and corresponding ChatGPT-developer conversation links. Each instance is labeled with a unique identifier: PN-X (X being the instance number) for patches not applied, NE-X for cases where no patches were suggested, and CL-X for closed pull requests. These identifiers support cross-referencing with the replication package.
@@
5.1 Results and Discussion – RQ2
Why are ChatGPT-suggested patches in conversations not directly integrated (PN) into the pull request, and how do developers use ChatGPT’s suggestions in their workflow under the self-admitted ChatGPT usage? In our analysis, PatchTrack initially classified 63 instances as patch not applied (PN). To validate the accuracy of this classification, we manually reviewed a sample of 55 PN instances, selected using a 95% confidence level with a 5% margin of error, as detailed in Section 3.2.3. During this review, we identified 3 false positives that were incorrectly labeled as PN and 4 false negatives from the patch applied (PA) set that should have been categorized as PN. After applying these corrections, we obtained a final set of 56 validated PN instances for thematic analysis. The analysis of these 56 instances revealed six distinct themes, as illustrated in Figure 6.
@@ 5.1 Results and Discussion – RQ2
Technical limitations and preferences: This was the second most prominent theme, with 9 of 56 instances demonstrating how ChatGPT’s patches were not applied due to technical constraints or developer preferences for alternative solutions. In these cases, developers faced trade-offs in implementation strategies, either due to limitations in the existing system or subjective choices regarding the best approach to a problem. A representative example is pull request #171 [63] within the ory/elements project (PN-11), where developers attempted to disable a button when submitting the form. A reviewer sought ChatGPT’s input on a possible fix, asking: “Wouldn’t the fix just be to change click to submit? […]” However, another reviewer pointed out a critical issue with this approach: “This doesn’t actually work. The reason is that the click event fires before the submit event […] but IMO we should just get this right […]” The pull request author ultimately acknowledged that resolving the issue properly would require additional time and effort, stating: “[…] just don’t want to spend time on getting this right, right now […]” Upon further examination, we found that the proposed solution was reverted because disabling the button caused unintended side effects, such as stopping event propagation. Given the minor impact of the issue, the team decided to focus on more pressing updates rather than implementing a workaround.
@@ 5.1.1 Implications for RQ2
While RQ1 revealed that developers rarely integrate ChatGPT patches verbatim, our RQ2 analysis adds a crucial nuance: even when patches are not applied (PN), ChatGPT influences developer decisions in meaningful ways. In these cases, developers assess suggestions critically and often extract conceptual guidance, architectural direction, or documentation support. This extends prior work [10, 11], which focused on patch acceptance rates, by showing that AI-generated code can shape developer thinking and workflow even when it is ultimately not merged.
Our contribution also complements recent studies examining the limitations of generative AI in software engineering. While prior work has investigated correctness risks in AI-generated code [37], cross-tool variability in patch quality [41], and the use of AI-generated metadata in PRs [20], our findings offer a behavioral lens into how developers evaluate and selectively reuse AI suggestions. Additionally, we build on research into multi-hunk patch challenges [38] and automation limitations in code review [36] by illustrating the human-in-the-loop decision-making that occurs when AI-generated patches are assessed but not applied.
Recent evaluations have highlighted limitations in LLMs’ ability to reason about and adapt to project-specific constraints [68, 69, 70]. These limitations likely contribute to patch rejection or modification. Our findings emphasize the need for AI systems that are better aligned with project context, coding conventions, and architectural patterns, an area that remains underexplored in current tooling.
@@ 5.2 Results & Discussion – RQ3
As shown in Table 2, PatchTrack classified 106 pull request instances as NE. For our analysis, we manually validated and examined a representative sample of 84 instances, as described in Section 3.2.3. From the generated summaries of the 84 NE instances, the employed card sorting analysis yielded four distinct themes, as illustrated in Figure 7. The identified themes are: Conceptual Guidance & Theoretical Advice, Documentation, Communication & Translation, Education and Knowledge Sharing, and Debugging & Optimization Strategies.
@@
These results build on and extend prior research on AI-assisted software development by highlighting a distinct usage pattern: developers often consult ChatGPT not for code generation but for conceptual refinement, documentation improvement, and knowledge support. While previous work has emphasized correctness limitations and patch quality issues in AI-generated code [37, 41], as well as the role of large language models in review automation [36], our findings point to a complementary function. They align with emerging views of large language models as collaborative assistants that support design thinking, aid code comprehension, and enhance communication, even in the absence of generated patches. These non-code uses suggest that ChatGPT contributes to software development in ways that extend beyond generation, offering new forms of value in pull request workflows that remain underexplored in prior literature.
@@ 5.2.1 Implications for RQ3
5.3 Results & Discussion – RQ4
How does ChatGPT influence closed pull requests, and what factors contribute to these outcomes under the self-admitted ChatGPT usage?
We identified 53 instances of closed pull requests (CL) for analysis. Following the same sampling methodology used in RQ2 and RQ3, we applied a 95% confidence level with a 5% margin of error, resulting in a stratified sample of 47 CL instances for manual inspection. We then performed an in-depth qualitative analysis of these 47 cases, focusing on the interactions between developers and ChatGPT within the context of the pull request discussions. The qualitative analysis identified six key themes, as illustrated in Figure 8. The most prevalent reason for pull request closure was attributed to Scope and Quality Issues, which accounted for 31.9% of the cases. Other reasons included procedural and technical adjustments (19.1%), experimental solutions (17.0%), administrative or policy issues (10.6%), duplicate solutions (10.6%) and cases where no specific reason (10.6%) was documented.
@@ 5.3 Results & Discussion – RQ4
Experimental Solutions: 8 of the 47 closed pull requests were later replaced by more refined implementations. These pull requests were often experimental, serving as proof-of-concept solutions that were ultimately superseded by optimized alternatives. In some cases, developers identified existing solutions that made ChatGPT-assisted contributions unnecessary for direct integration. For example, pull requests #2214 [77] and #2212 [78] in the open-learning-exchange/myplanet repository (CL-29 and CL-36) introduced ChatGPT-generated French and Arabic translations for UI strings, addressing issues #2216666https://github.com/open-learning-exchange/myplanet/issues/2216 and #2219777https://github.com/open-learning-exchange/myplanet/issues/2219. However, both pull requests were later closed, with their contributions consolidated into pull request #2229888https://github.com/open-learning-exchange/myplanet/pull/2229, which was merged. This closure pattern aligns with the common practice of refining and merging multiple contributions into a single comprehensive pull request. Although ChatGPT facilitated the initial translation drafts, final refinements were likely developer-led to ensure linguistic accuracy and adherence to project standards. This underscores how ChatGPT-assisted contributions often act as stepping stones toward more polished, project-aligned solutions rather than directly leading to merged changes.
@@
Prior studies have examined why pull requests are closed, highlighting factors such as contributor unfamiliarity, low code quality, and project maintenance policies [16, 17, 18]. Our findings complement this work by analyzing closures in the context of self-admitted ChatGPT usage. While some traditional reasons persist (e.g., scope misalignment, procedural issues), we identify new dynamics introduced by LLMs, such as AI-generated proofs-of-concept that are later superseded, or optimizations that, though valid, conflict with project-specific architectural norms. These patterns reveal emerging challenges in human–AI collaboration not addressed in earlier PR closure studies, and the persistent factors we observe align with established findings that PRs are often closed due to low code quality, misalignment with project goals, or procedural requirements such as policy compliance and workflow constraints [16, 17, 18].
@@ 5.4 Implications for Educators
6 Overall Findings and Discussion
Our study analyzed 338 pull requests across four distinct interaction categories: Patch Applied (PA), Patch Not Applied (PN), No Patch Generated (NE), and Closed (CL), to understand how ChatGPT influences collaborative software development workflows.
1. AI-generated patches are rarely adopted verbatim. In PA cases, the median integration rate was 25%. Qualitative analysis of 89 validated cases showed that developers most often engaged in structural integration, selective extraction, or iterative refinement, treating ChatGPT’s code as a starting point to be reshaped for style, maintainability, and architecture. High-integration cases, where patches were accepted with minimal edits, were rare, underscoring the need for more context-aware AI support.
2. ChatGPT adds value even when patches are not applied. In PN cases, developers used ChatGPT as a conceptual starting point, adapting or rejecting its code based on project constraints and maintainability goals. Its suggestions guided architectural and functionality decisions, validated alternatives, refined documentation, and improved project organization, influencing workflows even without direct code integration.
3. Non-code contributions shape development. In NE cases, ChatGPT’s responses influenced debugging strategies, naming conventions, design decisions, documentation quality, and knowledge sharing. These contributions highlight its utility beyond patch generation, supporting conceptual refinement, clearer communication, and informed decision-making.
4. Pull request closures reflect both traditional and AI-specific barriers. Consistent with prior findings on PR rejections [16, 17, 18], many CL cases were closed for persistent reasons such as low code quality, scope misalignment, or procedural or policy non-compliance. Others reflected emerging AI-specific patterns such as ChatGPT-generated proofs-of-concept later superseded, or valid optimizations conflicting with architectural constraints, highlighting new dynamics in human–AI collaboration.
5. Developer judgment remains central. Across all categories, developers critically assessed ChatGPT’s output against project requirements, coding standards, and maintainability goals. Suggestions were treated as one input among many, with final decisions shaped by architectural constraints, reviewer feedback, and domain expertise. This underscores ChatGPT’s role as a collaborator whose effectiveness depends on informed human oversight rather than autonomous decision-making.
These findings have broad implications for tool builders, researchers, and educators. Developers benefit most from AI tools when they can validate, adapt, or discard suggestions based on project-specific knowledge and constraints. Researchers should explore how to enhance AI’s awareness of repository context, architectural patterns, and team norms to support more effective reuse. Educators should equip students with critical thinking and review skills that emphasize human oversight in AI-assisted development. Together, these implications support a view of ChatGPT as a collaborator whose success depends on the judgment and engagement of the human developer.
@@ 7 Related Work
Pull request decision-making: In collaborative software engineering and open-source projects, the decision-making process for pull requests is crucial to determine both the quality and the efficiency of the development. Gousios et al. [15] and Yu et al. [82] have examined how contributions are evaluated, highlighting the importance of peer review and social interactions. Various factors influence pull request acceptance, including technical merit, code quality, and alignment with project goals. For example, Tsay et al. [16] found that reviewers prioritize maintainability, readability, and adherence to coding guidelines when evaluating contributions. Similarly, Zhao et al. [22] showed that pull requests that lack comprehensive tests are more likely to be rejected, reinforcing the importance of thorough validation. Complementing these findings, Soares et al. [18] identified interaction history and test inclusion as strong predictors of pull request acceptance, laying a baseline for comparison with emerging AI-assisted workflows.
Beyond technical factors, social and ecosystem-wide dynamics also influence pull request decision-making. Dey and Mockus [24] examined how technical attributes and the reputation of contributors affect the acceptance of pull requests in the NPM ecosystem, highlighting the importance of project constraints and community participation. Storey et al. [14] further emphasized that communication dynamics across different channels profoundly shape participatory cultures in software development, underscoring the collaborative challenges that arise in distributed workflows such as pull requests. In addition to social factors, prior research has highlighted the value of mining developer disclosures to better understand development practices. Yu et al. [83] introduced Jitterbug, a machine learning-based approach for detecting self-admitted technical debt (SATD) in code comments. Inspired by this idea, our study extends the concept of self-admitted artifacts to AI-assisted development by analyzing self-admitted ChatGPT usage in PR discussions.
@@
Beyond studies on AI-generated code, prior work has investigated how automation influences developer workflows and decision-making. Tufano et al.[36] found that automated code review tools, including those powered by LLMs, often ignore the human-in-the-loop context, limiting their reliability and developer trust. Cassee et al.[89, 90] showed that automated systems such as Continuous Integration (CI) and GitHub bots alter code review behaviors and are critically evaluated by developers based on trust and perceived utility. Extending this line of inquiry, Shihab et al.[91, 92] analyzed the adoption of LLM-based chatbots and automation bots in pull request workflows, emphasizing challenges related to trust, accuracy, and context sensitivity. Consistent with these concerns, Tanzil et al.[37] proposed a technique for detecting incorrect AI-generated code suggestions during code reviews, further underscoring the risks of hallucinated outputs.
In addition to studies on AI-generated code retention and quality, researchers have broadly examined AI-assisted software development, including the roles of ChatGPT and GitHub Copilot in code reviews, debugging, and software documentation [29, 30, 31, 93, 94, 95, 7]. Although these studies demonstrate improvements in productivity and code quality, they often focus on controlled experiments rather than real-world open-source development [32, 33, 34]. Recent work by Xiao et al.[20] analyzed how developers edit and adopt AI-generated PR descriptions, identifying frequent partial reuse and frequent human intervention – patterns echoed in our analysis of ChatGPT-suggested patches. Previous work has explored the role of AI in pull request workflow automation[96, 97], but little research has examined how developers interact with ChatGPT in pull request discussions and adapt AI-generated patches to fit project-specific constraints. Our work addresses this gap by empirically analyzing ChatGPT’s impact in pull request workflows, particularly the decision-making processes that influence patch acceptance, modification, or rejection. This challenge is further complicated when dealing with multi-hunk patches, which often require semantic coordination and broader context management. Nashid et al. [38] recently showed that LLMs struggle with such coordination across hunk boundaries, a limitation also observed in our dataset.